### v1

# Developers’ Perspectives on Software Licensing: Current Practices, Challenges, and Tools

Abstract.
Most modern software products incorporate open-source components, requiring development teams to maintain compliance with each component’s licenses. Noncompliance can lead to significant financial, legal, and reputational repercussions. While some organizations may seek advice from legal practitioners to assist with licensing tasks, developers still play a key role in such a process. To this end, it is essential to understand how developers approach license compliance tasks, the challenges they encounter, and the tools that they use. This work studies these aspects of software licensing practices through a study—conducted by a joint team of software engineering and legal researchers—consisting of a survey with 58 software developers and seven follow-up interviews. The study resulted in ???? key findings regarding the current state of practice. We discuss the implications of our findings and offer directions for future research as well as actionable recommendations for licensing tools.
1. Introduction
Software reuse, particularly through open source software (OSS), has become fundamental in the process of software development, allowing software developers to avoid duplicating efforts in areas where previous developers have found suitable solutions. However, the fact that a component is open source does not mean that it can be used without restrictions. Because software is protected by copyright and patent law (cop, 2021), OSS components are typically released under one or more OSS licenses that govern how a component may be reused, modified, and distributed. Failure to comply with the terms of any applicable license risks liability for infringement, which can result in legal, financial, and reputational consequences (Claburn, 2022; Vincent, 2022; Claburn, 2023; Gunningham et al., 2004; Vance, 2010; Gross, 2007), both for individual developers and for the organizations for which they work.
Ensuring license compliance throughout the development process is, therefore, a critical task. However, this process can be challenging to manage. Software systems can involve hundreds or thousands of OSS components, each under a separate license that must be identified and understood (Ballhausen, 2019). Furthermore, some OSS licenses have conflicting terms, making the use of components under different licenses problematic. There are hundreds of potentially applicable OSS licenses (spd, 2024; osi, 2024), some of which use legal terminology that is unfamiliar to developers and that does not always have a broadly accepted interpretation (Almeida et al., 2019; Gangadharan et al., 2012; Vendome et al., 2018).
As a result, it may seem appropriate to offload the work of license compliance to legal teams or Open Source Program Offices (OSPOs), but this may delay compliance efforts until near the end of a development cycle (Wintersgill et al., 2024), leading to lower productivity and wasted effort. Thus, while legal experts play an important role in compliance processes, it is also important for developers to be trained in such tasks.
As the individuals closest to the code, developers bear responsibility for maintaining compliance with the licenses governing that code (Wintersgill et al., 2024). However, where previous work has identified that licensing tasks can be challenging (Almeida et al., 2019, 2017; Vendome et al., 2018; Moraes et al., 2021; German and Hassan, 2009; Gangadharan et al., 2012), it does not identify developers’ perceptions of different licensing tasks, and questions remain regarding developers’ adoption of licensing tools and the strategies employed when addressing licensing issues.
To fill this knowledge gap, this study surveyed a variety of software developers, including those who may not perform compliance tasks regularly. Our goal was to gain a deeper understanding of the licensing processes and challenges for developers, discover the gaps in the technology developers use for licensing, and work toward organizational and technical solutions.
The main contributions of this paper are the following:
  • A novel assessment of how developers not only understand but engage with software licensing tasks, based on a survey of 58 developers with various levels of experience in licensing;
  • An in-depth analysis of the processes, roles, and tools that contribute to addressing software licensing issues, including the adoption of available tools and availability of resources in different contexts; and
  • Actionable insights into where future licensing tooling development should be directed, including key pain points and opportunities for generative AI.
2. Background and Related Work
2.1. Backgrounds on Software licensing
As software is protected by copyright law (among other legal doctrines) in the United States and around the world (cop, 2021), the owner of a piece of software’s copyright holds the rights to distribute, reproduce, or create derivative works of that software, as well as the right to authorize others to do so (USc, 2021). Such authorization often occurs by means of a license, which sets forth the permissible uses of the software and includes any restrictions on use. (Meeker, 2017) Use outside of the scope of the license constitutes copyright infringement, unless such use constitutes a fair use under U.S. copyright law.
OSS developers often choose one of several popular existing licenses to apply to their software, such as the MIT (mit, [n. d.]) license or a member of the GPL (gpl, [n. d.]) family of licenses. Permissive licenses, as the name suggests, typically allow broad use of the software, requiring only that, for example, the user provide notice files with information on the license or on attribution. Copyleft licenses, by contrast, constrain future use of the software by, typically, requiring that the user make the full source code of their work available under the same license. A developer must therefore understand, at a minimum, which licenses apply to components they incorporate into their work and what those licenses require.
Although several licenses have been widely used for many years, interpretive questions still exist. Besides the preliminary questions about the types of uses a license permits or prohibits, questions can arise about whether two licenses are compatible with each other, as well as how certain terms apply to desired software uses. For example, the requirements of GPL v2 apply to any work “based on the Program,” which the license defines as equivalent to “any derivative work under copyright law.” But this language does not answer the question of whether system calls to a library create a work “based on the Program” or whether the answer depends on whether the link is static or dynamic, a question that even OSS lawyers have identified as unresolved  (Wintersgill et al., 2024).
2.2. Developers and licensing
Previous work has shown that software developers often find licensing tasks difficult, such as determining whether two licenses are compatible with each other (Almeida et al., 2019, 2017), ensuring that their software complies with all relevant licensing requirements (Vendome et al., 2018), and releasing their software under multiple licenses (Moraes et al., 2021). Building on this work, we seek to contextualize these challenges by examining how often they are encountered and how developers address them.
Previous work also shows us that several methods exist to assist with licensing tasks. Both open source and proprietary tools can help with license compliance, such as Fossology (Fos, [n. d.]) and Black Duck (Bla, [n. d.]), respectively. Software bills of materials (SBOMs) can provide insights into software’s composition, facilitating licensing tasks (Bi et al., 2024; Stalnaker et al., 2024). Additionally, researchers have developed a variety of tools and processes aimed at assisting developers with compliance tasks (Tuunanen et al., 2009; Ombredanne, 2020; Kapitsaki et al., 2017; German and Hassan, 2009; German and Di Penta, 2012; German et al., 2010; Hemel et al., 2011; Feng et al., 2019; Gangadharan et al., 2008). Some of these studies have explored the use of generative AI to assist in compliance tasks (Cui et al., 2025; Kahol et al., 2025; Ke et al., 2025); Other studies, based on mined software repositories and Q&A websites, have explored methods to detect non-approved licenses, non-standard license variants, and exceptions.  (Meloca et al., 2018; Zacchiroli, 2022; Papoutsoglou et al., 2022; Vendome et al., 2017b; Xu et al., 2023). Where previous work has sought to uncover ways to address licensing issues, our work seeks to understand if and how developers are utilizing these methods.
Prior work has also explored some of the processes and decisions developers make during license compliance efforts. Wu et al. conducted a large empirical study aimed at understanding license usage (Wu et al., 2024). Works by Vendome et al. and Di Penta et al. sought to understand when, how, and why developers opt to change the license of their software (Di Penta et al., 2010; Vendome et al., 2017a). Liu et al. introduced a method to predict license changes from source code changes (Liu et al., 2019). Relatedly, Wintersgill et al. surveyed and interviewed legal professionals with experience in software license compliance to understand their challenges and their views on software license compliance tasks (Wintersgill et al., 2024). Lastly, Li et al. provide a systematic literature review of stakeholders’ challenges in license compliance tasks (Li et al., 2025).
Where the literature offers insights, solutions, and tools for licensing tasks, this work seeks to expand on previous efforts in two key ways: (1) by understanding if and how developers adopt these tools and techniques in practice, and (2) by identifying the license compliance pain points that are most relevant to developers by providing greater insight into the specific problems developers report encountering most often, which they report are most difficult, and what strategies they use to address them.
3. Study Methodology
The goal of this study is to analyze the processes employed and challenges faced by software developers in performing OSS license compliance tasks. The study was conducted through a survey and follow-up interviews, and involved a collaboration between software engineering (SE) and law researchers. The context of the study consists of 58 developers recruited through personal contacts and open-source projects’ discussion channels.
More specifically, the study aims to address the following three research questions (RQs):
  1. • RQ1:
    What are the licensing challenges faced by software developers in practice? This RQ explores the licensing compliance challenges developers face in their routine workflows.
  2. • RQ2:
    What are the processes by which software licensing problems are resolved? This RQ complements the previous one by exploring how developers resolve compliance challenges when they occur, and investigating how developers perceive the relationship between development and legal teams.
  3. • RQ3:
    How are developers incorporating license compliance tools in their workflows, if at all, and what issues have they encountered? Here we explore how licensing tools are being used by developers, if at all, as well as what issues and shortcomings developers report when using such tools.
This work, including the survey questionnaire, the participant identification procedure, and the survey and interview protocols, was approved by our institution’s ethics review board. An overview of our methodology is depicted in Figure 1.
Figure 1. Research methodology (image credits at (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2025))
3.1. Survey design and participant identification
When designing our survey questionnaire, we considered previous literature on software license compliance, as well as general guidelines (Groves et al., 2009) and SE-specific guidelines for survey design (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). The survey went through multiple review cycles, involving seven SE researchers and one law researcher, to achieve appropriate content, clear and concise language, and eliminate confusion and bias. We sought to limit the response time of the survey to minimize the cognitive load on participants and to encourage higher completion rates. In addition, we conducted a small pilot study with members of our research laboratory to further refine the survey questions, including by identifying questions that could be culled or reworked to better focus on the study’s research goals. The final survey included questions to obtain information related to each of our RQs, including questions relating to the compliance challenges faced by developers, existing compliance practices, and the use of tooling or other automation to perform compliance. Figure 2 provides an overview of our survey questionnaire. The complete survey text can be found in our online replication package (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2025).
Figure 2. Survey questionnaire outline
To recruit developers with diverse experiences, we distributed the survey in three ways: (1) through our professional networks (e.g., posts on LinkedIn and emails), (2) through snowball sampling, encouraging respondents to share the survey within their networks, and (3) by posting links to the survey on relevant servers on Discord.
To identify Discord communities, we examined the TIOBE index (Jansen, 2025) of top programming languages to identify the top 20 most popular programming languages as of January 2025. We then searched for Discord servers associated with these languages (excluding Scratch, which is primarily a tool for beginners) by locating server links through the languages’ own documentation or websites, through third-party Discord server browsing tools DiscordMe (DiscordMe, [n. d.]) and Disboard (Disboard, [n. d.]), and through Reddit posts on subreddits related to the targeted languages. This process was further supplemented by adding three generalized programming/development-focused servers and one server each for the popular React and Node.js frameworks. Altogether, this yielded 27 Discord servers representing 16 of the top programming languages as well as general programming spaces and popular frameworks. Next, one researcher carefully analyzed the rules of each Discord server to determine whether it would permit a solicitation for participation in a survey, asking server administrators in cases where an answer was not immediately clear. In total, we were able to post survey invitations in 14 software development-focused Discord servers, with total membership exceeding 485k users at the time of distribution. Two additional posts were made in smaller servers but were later removed by moderators.
3.2. Survey response collection and analysis
Survey responses were collected using Qualtrics (qua, nd) from February 2025 to June 2025. Survey invitations were sent out iteratively during this same period. To monitor where responses originated from, we distributed survey URLs with identifiers unique to each source of participants. We obtained 58 complete survey responses. Of these, 38 came from our professional networks and 17 came from Discord, while 3 participants seemingly removed the tracking component from the survey URL. The largest population from a single Discord server was 6 participants.
Responses to the five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Three authors, all SE researchers, performed open-coding by independently assigning one or more codes to each response using a shared spreadsheet and codebook. Each annotator independently coded all 58 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes. Reasons for disagreement included different interpretations of potentially ambiguous responses, applying the same codes in different contexts, and using broader codes rather than more specific ones. All such disagreements were resolved among the annotators, resulting in a final set of codes, definitions, and assignments for each response. Our replication package (Nathan Wintersgill, Trevor Stalnaker, Daniel Otten, Laura A. Heymann, Oscar Chaparro, Massimiliano Di Penta, Daniel M. German, Denys Poshyvanyk, 2025) contains our final codes and definitions. We did not base our analysis on inter-rater agreements, as the nature of the study was such that codes were developed through an inductive process, and multiple codes could be assigned to each survey response. However, we carefully followed best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the reliability of the results.
We have corrected grammar and spelling errors in quotations from open-ended questions for readability and have indicated by ellipses or brackets any deletions or additions to text for clarity or space-related reasons. All responses have been attributed to survey participants using anonymous IDs for traceability (e.g., $R_{12}$).
3.3. Interview design and analysis
To better understand participants’ responses and gain further insights into their views on software licensing and related issues, we conducted semi-structured follow-up interviews over Zoom. At the conclusion of the survey, we provided respondents with an option to leave an email address if they would be willing to participate in a follow-up interview, which 21 participants utilized. We contacted all of these individuals, which led to seven respondents agreeing and participating in follow-up interviews.
Each interview lasted 20 to 30 minutes and was recorded to facilitate transcription and subsequent analysis. The interviews were transcribed using Whisper (Whisper, [n. d.]) and reviewed against the original recordings for accuracy by one researcher each. Three reviewers assigned topic labels to relevant portions of each transcript, and all three reviewers then discussed the labels to reach a consensus.
To preserve interviewee anonymity, we have assigned each one an identifier to attribute the source of quotations (e.g., $RI_{4}$). As with the survey results, we have corrected errors in grammar for readability and have indicated deletions or additions by ellipses or brackets.
3.4. Participant demographics
Respondents reported a variety of different backgrounds and experiences. Twenty-one countries were represented, with the largest groups of participants residing in the United States (22, 38%) , Germany (6, 10%), Bangladesh (5, 9%), and Colombia (5, 9%). Forty respondents (69%) reported that they worked on proprietary software, 33 (57%) on open-source projects that relied on third-party components, and 14 (24%) on open-source projects with no dependencies. The presence of these different groups bolsters our results by allowing us to capture information on licensing tasks for different use cases. Most of the respondents described their primary role as programmer (32, 55%), with other common roles including researcher (10, 17%), project technical lead (5, 9%), tester (3, 5%), educator (3, 5%), and software architect (2, 3%). Development experience ranged from less than one year to 35 years, with an average of 9.9 years and a median of 7 years. Domains also varied broadly, ranging from contracted government work and healthcare to game development and banking. Almost two-thirds of the respondents had experience with web development (37, 64%). Other common system types developed by the respondents included libraries/frameworks (25, 43%), desktop applications (20, 34%), mobile applications (15, 26%), development tools (14, 24%), middleware (13, 22%), and AI-intensive systems (9, 16%). Although not extensive, the developers in the response pool represent a broad range of (self-reported) experiences and backgrounds, with likely diverse experiences with the software licensing process.
We evaluated participants’ experience with licensing by asking them if they had any training or education in software licensing, and if so, whether that training was formal (e.g., college classes, degree, or certification) or informal (self-learning, employer-provided, etc.). As shown in Table 1, more than half of the respondents indicated that they had some kind of training (35, 60%), although most reported that it was informal training (25, 43%). We note that our goal was to investigate the perspectives of developers broadly: while licensing experts and compliance specialists can be expected to know more about licensing issues and may have different strategies for addressing them than an average developer, everyday developers are on the front lines of license compliance (Wintersgill et al., 2024). Capturing their perspectives grants insight into how compliance is actually done (or not) by the people who encounter the need for it regularly.
4. Results
In the following, we present the results of our study, which emerged from the survey responses and interviews collected.
4.1. RQ1: Developer Licensing Challenges
Training Count Percentage
No 23 39.66%
Yes, informal 25 43.10%
Yes, formal 4 6.90%
Yes, formal and informal 6 10.34%
Table 1. Respondents’ license training
This section discusses participants’ responses to questions regarding the different software licensing tasks they perform and the challenges they face.
4.1.1. Most common challenges
Participants were presented with a list of software licensing-related tasks and asked to identify how frequently they engaged in each task on a scale of “never,” “occasionally,” “frequently,” or “all the time.” The tasks were derived from our previous experience working with licensing issues and from prior work cataloging licensing bugs (Vendome et al., 2018). As shown in Figure 3, the tasks developers reported most often engaging in (i.e., frequently, or all the time) were identifying the licenses of software components they depend on, understanding the terms of use/service associated with the tools they use to develop software, and verifying that their software complies with all relevant licenses. Developers reported that they assigned a license to their own software slightly less often, though this was the task most reported by the respondents as being performed “all the time” (10). By contrast, most respondents indicated that they never spent time changing their software’s license (42) or acquiring or granting license exceptions (40), and slightly less than half reported that they never spent time identifying key terms from software licenses (28). Notably, developers reported performing all tasks “never” or “occasionally” more than they reported performing them “frequently” or “all the time,” suggesting that they perform licensing tasks relatively infrequently overall.
Figure 3. How frequently developers perform licensing tasks
Finding 1: While developers broadly reported that licensing tasks may not be common, the ones they engage with most are identifying their dependencies’ licenses and understanding the license terms associated with development tools, as well as verifying that their software is in compliance with all relevant licenses and assigning a license to software.
Changing a software project’s license can be difficult (Wintersgill et al., 2024), particularly for open-source projects, but such changes do occur (Vendome et al., 2017a). Few developers, however, reported having experience with changing the license of a software component, and no respondents reported performing this task “all the time.” Similarly, most developers reported that they had no experience in acquiring or granting license exceptions, in which an exception to the terms of a component’s license is granted to a particular licensee (for example, to grant additional permissions or remove restrictions that would otherwise be in place). License exceptions can be difficult to obtain from open source projects (Wintersgill et al., 2024) and may be a process with which developers are unfamiliar; indeed, this category received the most “I don’t know” responses, as shown in Table 2. Almost half of the respondents reported that they never spent time identifying key terms from licenses. This may suggest that developers spend little time analyzing the contents of software licenses, or that they feel they understand available software licenses and how they interact, so they do not need to spend further time reviewing them.
Finding 2: Most developers reported never changing their software’s license or obtaining/granting license exceptions. Almost half of developers never dedicate time to identify the key terms of software licenses.
4.1.2. Most difficult challenges
Developers were also asked to rate the difficulty of licensing tasks using a 5-point Likert scale from “extremely difficult” to “extremely easy.” Here, as shown in Figure 4, we observe that tasks such as identifying software licenses, understanding terms associated with development tools, and assigning a license were viewed as easiest, while changing a license, dealing with license exceptions, and identifying key license terms received more mixed responses. Notably, respondents did not reach a consensus for any task, showing that it was easy: the results show mixed opinions at best. Our results confirm previous work showing that developers can find licensing tasks challenging (Almeida et al., 2017, 2019).
Finding 3: Developers gave mixed answers regarding the difficulty of most licensing tasks examined. No tasks were broadly considered easy.
Figure 4. Developers’ perceived difficulty of licensing tasks
Verifying whether software complies with all relevant licenses was cited as the most difficult task by respondents, being the task most frequently labeled “somewhat difficult” (25) as well as most frequently labeled “extremely difficult” (13). This is particularly notable, as compliance verification was reported as one of the most frequently performed tasks in the previous question. Given that this task can occur whenever components are added or changed, it is possible that its perceived difficulty dissuades developers from performing it. Identifying key license terms was also frequently cited as somewhat/extremely difficult. That developers struggle to understand licenses and their terms, and that verifying compliance proves a difficult task for them, is a widely accepted truth in this area (Almeida et al., 2019; Gangadharan et al., 2008, 2012; Vendome et al., 2018). Our results confirm these previous findings while providing additional depth by demonstrating which tasks developers find most challenging among these, and contextualizing them with details on how often these challenging issues arise during development. This further motivates the need for research efforts to improve the process of confirming software compliance.
Finding 4: Verifying license compliance, while one of the most common licensing tasks, was also reported to be the most difficult. Identifying key license terms was also frequently perceived as “extremely difficult.”
We also note that a sizable group of respondents answered “I don’t know/no opinion” to the survey questions regarding licensing task frequency and difficulty, as shown in Table 2. In particular, we highlight the large number of “I don’t know” answers regarding the difficulty of changing a license (30) or receiving/granting license exceptions (34), corresponding with similarly high numbers of respondents who reported that they never performed these tasks and indicating a lack of familiarity with them among developers.
Task How often How difficult
Identifying dependency licenses 0 7
Understanding dev. tool terms 1 6
Assigning a license 1 14
Changing a license 2 30
Verifying compliance 2 12
License exceptions 4 34
Identifying key license terms 3 17
Table 2. “I don’t know” / “I have no opinion” responses to task frequency and difficulty questions
4.1.3. Licensing factors
To better understand the elements involved in the decision to license software, participants were asked to identify the factors they or their organizations consider when assigning licenses to software. Participants could select multiple such factors. As shown in Figure 5, developers most commonly cited legal considerations surrounding their software (30) as influencing the license assigned to their software. Together with the second most popular answer, organizational policies (24), this illustrates how licensing decisions are often made for non-technical reasons. The third most frequent answer, dependency licenses (23), shows how constraints based on third-party components can similarly affect such decisions. Only beyond this point do we observe specific considerations related directly to development-focused concerns, such as the intended distribution of the software project (21) or its intended use (19). Philosophical considerations, such as the participant’s or their organization’s views on ethical software distribution and use (13) or their views on the nature or purpose of OSS development (12), received fewer responses. Overall, these results highlight a hierarchy of considerations when choosing a license: first, any legal obligations or organizational policies must be adhered to, then technical limitations based on the project and how it should be used, followed by philosophies regarding software and its reuse. It should be noted that as participants were able to select multiple responses, a given respondent could highlight multiple of these groupings to be considered in tandem with each other, and for some respondents, the categories may overlap. However, the preeminence of non-engineering or non-usage requirements illustrates the first priorities respondents reported having when making licensing decisions. As such policies may vary among jurisdictions or among organizations, we suggest that licensing tools should accommodate organizational choices.
Finding 5: When considering how to license software, the surveyed developers appeared to first consider legal and organizational obligations before applying technical constraints based on the project and then, less commonly, their philosophy on licensing or OSS.
Figure 5. Factors considered when assigning a license
4.2. RQ2: Processes for Addressing Licensing Problems
Here, we discuss how the study participants reported resolving licensing problems, including the involved roles, the adopted resolution process, and how participants understand licenses.
4.2.1. Roles
Legal professionals often participate in licensing compliance tasks, either in-house or as outside counsel (Wintersgill et al., 2024). We asked our participants whether they or their organization regularly worked with legal professionals or people otherwise trained in software licensing and/or copyright law. Eighteen participants responded “Yes,” 25 responded “No,” and 15 responded “I don’t know.” This indicates that expert-level licensing knowledge may not always be available to developers—or, in cases where it is, they may not be aware of it. This creates a particularly challenging situation since developers often do not have training associated with software licensing and find licensing tasks challenging (Almeida et al., 2019).
To understand developers’ perspectives on the issue in practice, we asked participants, “If an issue related to software licensing arises during software development, who is responsible for resolving that issue?” Participants could select developers, legal experts, an OSPO, or enter another role under “other.” As shown in Figure 6, respondents were divided on the issue. Although legal experts were most frequently cited as being responsible for addressing licensing issues (23), developers were also frequently listed (17). Notably, 5 respondents who answered that they did not regularly work with legal professionals and 4 respondents who reported not knowing if they or their organization regularly worked with legal professionals nonetheless selected legal experts as responsible for resolving licensing issues. Several respondents indicated they did not know whose responsibility it was (20), which may indicate a lack of experience or awareness in this topic. Developers rarely cited an OSPO, a group dedicated to managing activities related to open-source software, as responsible for addressing licensing issues (7). One respondent selected “other” and elaborated in a short answer that they were solely responsible for licensing in their case.
The number of respondents who viewed legal experts as responsible for addressing licensing issues (23) exceeds the number of developers who indicated that they regularly worked with such experts (18). This may indicate that the processes for addressing licensing issues are unstandardized and may not meet stakeholder expectations. We discuss the implications of this in Section 5.1.
Finding 6: Developers are divided on who is responsible for addressing licensing issues. Similarly sized groups chose legal experts and developers as responsible, while just as many did not know who was responsible.
Figure 6. Roles perceived by developers as responsible for addressing licensing issues
4.2.2. Problem resolution process
To gauge the difficulty that developers had when addressing licensing issues, we asked survey participants to answer whether solutions to licensing issues were typically immediately apparent or if such issues required research. Participants were invited to elaborate. Of the 33 respondents who indicated that they had experienced licensing issues, only 6 respondents (10.34%) reported that solutions were readily apparent. Twenty-one respondents (36.21%) indicated that addressing licensing issues required additional research, while 6 respondents (10.34%) said that it depended on the nature of the problem.
Many respondents then provided details on how they might address such problems. The most common such recourse was to involve a legal team in the process, as specified by 9 respondents (15.52%). Sometimes, this was represented as an interactive process involving both lawyers and developers: $R_{33}$ noted that “[t]he complexity depends on the type of license, usage rights, and legal implications, so it usually involves reviewing the specific terms of the license and possibly consulting legal or compliance experts.” In other cases, the process involved less “consulting” in favor of passing the issue to others, such as $R_{40}$, who said that issues went to “[m]y manager, prior [to contacting] the legal advisory team,” and $R_{38}$, who noted that “[f]urther investigation is passed to an expert in corporate law.” However, we note that such resources may not be available in all cases: for example, open-source development projects may have different resources available than corporate ones. Given that this was the most-reported strategy, this demonstrates a need for more accessible options for addressing licensing issues, or, at least, better adoption of such strategies. Four respondents (6.90%) stressed that an understanding of the involved software licenses and their terms or requirements is a necessary prerequisite to solving licensing problems. As $RI_{2}$ indicated in their survey response, to find solutions to licensing issues, “you need to fully understand the licensing in the first place.”
Finding 7: When faced with licensing issues, developers’ most-reported strategy was escalating the issue to legal experts.
The respondents also identified several specific strategies that could be deployed once a licensing issue was identified. Three respondents (5.17%) indicated that when issues with a dependency’s license arise, they might search for an alternative component with similar functionality. Two of these (3.45%) further noted that, as a last resort, they might write an internal alternative if a third-party solution were not feasible for licensing reasons. $R_{14}$ explains: “Usually we pivot to a different dependency with a compatible license or ask for license exceptions if possible. We had one case where we had to resort to doing an [in-house] development because the only library that matched the requirements was incompatible.”
$RI_{6}$ noted that, concerning the timing of licensing tasks in the software development cycle, “the bulk happens at the first release…” This aligns with prior work showing that many organizations treat licensing as something to be addressed at the end of development, although such an approach introduces new challenges (Wintersgill et al., 2024).
Some responses identified solutions that appeared only once (1.72%) in our results, including reverting to an older version of a dependency if that version had a more favorable license, purchasing a license when available, simply ceasing use of the affected component, or letting the issue go unaddressed. While the methods identified here align with previous work (Xu et al., 2023), our results show that no method was particularly favored, highlighting the need for licensing solutions and, by extension, tools that can offer solutions tailored to the specific circumstances of the software in question.
Finding 8: Developers presented many strategies for addressing licensing issues at the dependency level, including changing dependencies, reverting to an old version, or implementing functionality themselves.
4.2.3. Understanding licenses
In interviews, we sought to clarify how developers acquired an understanding of software licenses. $RI_{3}$, who had indicated that they had received informal training (see Table 1), described a “bootcamp” for licensing consisting of “three sessions of 45 minutes.” Some interviewees described how they had gained an understanding of common licenses over the years, such as $RI_{5}$, who noted, “I generally know that if it’s MIT or Apache, I’m good to go… and those are the ones that I encounter most regularly.” $RI_{6}$ had a similar familiarity with their most-used licenses but acknowledged that this did not extend to licenses they had less experience with: “I’m familiar with… the [Creative Commons] sharealike, the MIT license… But I wouldn’t say that I’m comfortable knowing what all the most common ones mean. I know the ones that I know because they are the ones that I use.”
Interviews also revealed how developers learned about unfamiliar licenses encountered during development. For example, $RI_{7}$ described “[searching] mainly on the web, and looking at other cases. And also the basic Wikipedia…. [I once used] ChatGPT… [for] interpreting the rules that are written [and] making sure that I understood well.” Developers may also look for key terms in software licenses that pertain to them, as explained by $RI_{5}$: “Mostly what I’m looking for is, what are the restrictions on how to use it, and then, what are the particular attribution requirements or restrictions for it.” The terms of interest can vary depending on the use case or organizational policies, such as in the case of $RI_{4}$, who noted that “as far as [my organization] goes, there is a strict policy for patent licensing and getting licenses for IP strategy.”
Finding 9: Developers report best understanding licenses they use most frequently, but may search online or use resources to understand other licenses, paying attention to terms that they believe most affect them.
4.3. RQ3: Licensing Tools
To understand if and how developers use tools to assist with software licensing tasks, we asked participants to describe their licensing tool use. Here, we refer to a “licensing tool” as any software that can assist with one or more software licensing tasks.
4.3.1. Tool adoption
Of 58 complete responses, only six participants indicated that they had used software tools to assist with software license compliance. This low adoption rate might suggest a lack of awareness or interest in licensing tools, despite the perceived difficulty of some licensing tasks, as discussed in Section 4.1.2.
Interview participants elaborated, suggesting that the frequency with which licensing tasks are performed may be a factor in low tool use. $RI_{5}$ stated, “I’m not aware of tools [because of] the fact that [licensing is] not something that I have to deal with every day…” Similarly, $RI_{6}$ was unfamiliar with tools, but mused that someone who engaged with licensing more might be more familiar: “I am less aware [of licensing tools]… [but] I would be completely unsurprised if there are established open source tools to assist in this [that] I am not personally aware of, but the people that work for me and do this when it is an important task very well may [know of them].”
Finding 10: Adoption of software licensing tools was low among respondents, who suggested that this may be due to the relatively low frequency at which licensing tasks are performed.
4.3.2. Tool benefits and shortcomings
We asked those participants who indicated that they used licensing tools about their experiences with such tools to better understand how they are used in practice. First, respondents were asked to identify the benefits these tools provided. Given the small size of this group, no clear trends emerged, although we can extract some notable reasons for using such tools. Participants expressed that tools were useful in identifying the set of licenses present within software and its dependencies, including $R_{4}$, who noted that tools were useful for “automat[ing] recursive gathering of licenses,” and $R_{43}$, who appreciated tools that could increase “[v]isibility on transitive dependencies” and “point out troubling dependencies.” $R_{1}$ went a step further, claiming that tools help them “ensure licenses are always compatible on every release.”
Respondents also indicated that tools could help them to choose licenses to assign to their software. In their survey response, $RI_{7}$ indicated that they benefited from tools that could give “an overview of all licensing options [and support] the selection of the right option.” They also noted a capability of “[g]enerating terms that I can copy/paste,” though it is not clear whether this referred to providing the text of existing licenses or creating new licenses. In interviews, $RI_{5}$ and $RI_{7}$ both referred to using choosealicense.com (GitHub, Inc., [n. d.]) to assist them when assigning licenses to software. Specifically, they appreciated its ability to “answer your questions and then recommend the best license,” as $RI_{7}$ put it.
Finding 11: Developers highlighted the capability of tools to assist with identifying licenses in dependencies, highlight those which may cause licensing issues, and choose appropriate licenses for their software as being particularly beneficial.
We also asked respondents who had used licensing tools to describe any particular difficulties or shortcomings that they had encountered. Two respondents indicated that they had encountered issues with licensing tools being cumbersome or difficult to use, with $R_{43}$ describing usability as “poor.” Others noted that tools were not able to account for user feedback: in their survey response, $RI_{7}$ called current tooling “[n]ot interactive enough,” and noted that it could be improved with an “[i]nteractive decision tree” or a “[r]eliable chat bot.” Other concerns included the clarity of tools’ outputs, such as $R_{43}$’s observation that the “[a]ctual effect of a license is obscure[d] to developers,” and $R_{37}$’s opinion that tools “[m]ay abstract away finer details.”
Finding 12: In the views of respondents, licensing tools are hindered by low usability, low interactivity, and a lack of clarity and depth to analysis.
4.3.3. Areas requiring better tool support
Given the low reported tool adoption, in interviews, we investigated which licensing tasks respondents believed could most benefit from better tool support.
In interviews, respondents noted that ways to better explain licenses and their terms would be beneficial. To illustrate this, $RI_{5}$ cited how “GitHub [has a] drop-down list when you choose a license, but it doesn’t tell you what [the licenses] mean.” $RI_{3}$ echoed this sentiment, elaborating that interpreting the terms of software licenses was a challenge: “[T]here is a gap between what the license gives you as $\dots$ content and [how] you can interpret it $\dots$ [You] can understand the text, but if you need to learn how to interpret that, that would be nice, a tool that gives you the interpretation.” The challenge of interpreting licenses intensifies when considering software that crosses borders and licenses that were written in other legal and cultural contexts, as identified by $RI_{7}$: “[W]hen you are reading the laws sometimes, you have to interpret and it’s very hard to make sure that your interpretation is the right one. And since we are talking in an international context, sometimes it depends $\dots$ there [are] some cultural aspects that make you interpret[ ] the law in [different] ways. But when you are talking about something which is worldwide, it’s very difficult to interpret how the people that designed the rules thought about it.”
Similarly, $RI_{7}$ sought a tool to highlight the differences between licenses: $\dots$ I would like to have something to compare licenses. Because I think that there are some overlaps between some licenses, and sometimes it’s hard to see the borders of the license.”
Beyond this, some interviewees expressed a need for tooling support for some of the more difficult compliance tasks, as defined in Figure 4. For example, $RI_{7}$ noted that they did not know of any tools for verifying license compliance.
Noting developers’ difficulties in understanding software licenses, $RI_{4}$ identified a need for assistance in communicating with legal experts: “[T]here [could] be a feature that can guide technical people like us, about the legal terminologies better $\dots$] as well as the communication between the legal people and the technical people. $\dots$] For example, what kind of things do we need to talk about?”
Finding 13: Respondents believed that licensing tools should support some of the more difficult licensing tasks identified by developers and assist developers with communicating with legal experts for those tasks.
[…]
Finding 14: Developers expressed mixed feelings on the potential of generative AI with respect to software licensing: the impacts it might have are not yet clear, and further study is required.
[…]
Overall, licensing tools with sufficient capabilities may improve the process of performing licensing tasks in various ways, as long as they are sufficiently reliable and accessible.
Finding 15: Respondents conveyed that software licensing tools should be reliable and easily accessible by developers, such as through existing tools and platforms.
5. Discussion
This section outlines the implications of our findings for practitioners and suggests areas that need additional focus.
5.1. Who is responsible for license compliance?
In Section 4.2.1, we observed that respondents were divided about who exactly should be responsible for handling license compliance tasks, but the most common answer was that legal experts should ultimately be responsible. Accordingly, in Section 4.2.2, we observed that when faced with licensing issues, respondents’ most-reported strategy was indeed to escalate to legal experts. However, we also observed that such experts are not always available to developers.
This result is in contrast to the findings of another recent work that considered how legal practitioners approach software license compliance challenges (Wintersgill et al., 2024), which found that legal professionals believed that compliance should be a team effort, but that the primary responsibility for compliance should fall on developers, who are most familiar with the project. Legal professionals may have limited technical training and are not closely involved in the code base. In their view, it was the role of legal teams to provide developers with the tools, resources, and training they need to succeed and to provide guidance on only the more challenging issues.
It would seem that both developers and legal practitioners believe that the ultimate responsibility for compliance should fall on the other party, and both provide valid arguments for why this should be the case. In truth, the answer likely lies somewhere in the middle, with an appropriate collaboration of legal and development teams that can bring the optimal mix of technical knowledge and legal background. However, as noted by Wintersgill et al. (Wintersgill et al., 2024) and Section 4.3.3, this balance can be difficult to strike, since legal experts and software developers may have different interests and goals. More research is needed to explore the interaction between roles and determine ways to facilitate effective collaboration.
5.2. Where to focus future efforts?
While our results show that licensing tasks may be infrequent for developers, their importance is undiminished: failing to comply with licenses can result in significant costs (Claburn, 2022; Vincent, 2022; Claburn, 2023; Gunningham et al., 2004; Vance, 2010; Gross, 2007), justifying focused investment in tools to assist with licensing tasks.
As shown in Section 4.1.1 and Section 4.1.2, developers perform software licensing tasks at different rates, and some tasks are more difficult than others. In particular, while verifying license compliance was reported as one of the development tasks developers perform most frequently, it was also cited as the most difficult task by far. Despite this, respondents were broadly unaware of tools capable of assisting with this task. Section 4.3.1 showed low tool adoption, mainly because of the lack of tooling better integrated into the software development process. As suggested in Section 4.3.3, licensing tools built into an IDE or into a commonly-used service like GitHub would make their capabilities easier to access for developers, while also mitigating usability concerns highlighted in Section 4.3.2. Increased accessibility may also mitigate the last minute timing of resolving licensing issues, as identified in Section 4.2.2.
The broad array of licensing factors considered (Section 4.1.3), roles involved (Section 4.2.1) and potential solutions to issues (Section 4.2.2) indicate that the licensing tools must be able to take into account the context provided by the user and adapt to the specific circumstances of a given project.
Section 4.3.3 also articulated the developers’ varying views on generative AI as it relates to software licensing. Our interviews revealed that while some are hopeful that AI-based tools could assist with software licensing in new, powerful ways, others expressed caution regarding their use for such a critical purpose. Therefore, we propose that an investigation into the suitability of generative AI for licensing tasks is warranted. An AI-based licensing tool would need to be capable of providing meaningful analyses of software licenses and software components’ licensing status, while also being reliable enough to avoid making inappropriate recommendations or providing services best left to legal practitioners.
6. Threats to Validity
Construct Validity threats concern the relationship between theory and observation. As with survey studies generally, a possible threat may be related to the fact that our study can only report participants’ perception, rather than actual practices. Interviews partially mitigated this threat as they allowed us to seek clarification and expansion of responses. However, different kinds of studies, e.g., ethnographic studies, would be required to better understand developers’ practices in the field.
Internal Validity threats concern factors internal to the study that could influence our findings. One possible threat is subjectiveness in coding interviews’ open-ended answers and interview transcripts. To mitigate such a bias, we utilized an open coding methodology with a procedure to discuss and resolve conflicts. We employed diverse strategies to locate participants (professional networks, Discord servers, etc.) in an attempt to increase the pool of different perspectives and minimize potential bias, but we are aware of the issues that may arise from low response rates and self-selection bias. We followed best practices in formulating survey and interview questions, making sure that questions were written clearly and concisely to avoid confusion and biased language and conducting a small pilot study. Interviews were limited to at most half an hour, meaning that not all of a participant’s views were likely heard. We limited confirmation bias in qualitative analysis by independent coding, discussing disagreements, and arriving at a consensus based on the data.
External Validity threats concern the generalizability of our findings. The conclusions drawn in this study apply only to the population that participated in our survey and follow-up interviews. Our results cannot be generalized to the larger population, but in light of the themes that emerged, we believe that other software developers likely share some of the same experiences and challenges. That said, our goal was not to claim generalizability but to identify potential pain points that hinder developers from achieving efficient and effective license compliance for their projects.
7. Conclusion
We present a comprehensive study of current software licensing challenges faced by developers through a survey of 58 developers and seven follow-up interviews, deriving ???? key findings encompassing: (i) how frequently different licensing tasks are performed; (ii) their difficulty; (iii) the roles developers perceive as addressing licensing issues, (iv) processes for resolving those issues; and (v) the benefits and shortcomings of licensing tools, and opportunities to improve them. These results reveal actionable insights for the creators of licensing tools by highlighting areas that need improvement, such as low perceived interactivity and shallow analyses, as well as identifying both common and difficult licensing tasks as areas that can most benefit from future license tool development.
Respondents had different strategies for handling licensing issues and differing views on who is responsible for addressing them. Although many reported performing licensing tasks infrequently and without licensing tools, they also highlighted difficult tasks, particularly verifying license compliance. As such, we recommend that future research and tool development focus on pain points and investigate the suitability of generative AI for licensing tasks.